OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)).

OpenStudy (anonymous):

f(g(x)) = -4x + 7 In other words, if g(x) = 10x -6, it can be f(10x-6) instead of f(g(x)). Now you replace it with your x.

OpenStudy (anonymous):

f(10x-6) = -4(10x-6) + 7 and solve.

OpenStudy (anonymous):

ok thanks that actually was helpful and i got my answer faster than i thought i would thank you

OpenStudy (anonymous):

No problem, any time ;)

OpenStudy (anonymous):

Please close your question if it has been answered. Thank you.

OpenStudy (goformit100):

Giving you an energetic and warm Welcome to one and only 'Open Study'.